OpenStudy (anonymous):

last semester i got GPA=3 ,if the GPA for the next semester is always the average of the previous one.Suppose i get GPA=4 for the next 8 semesters,is it possible to end up with average GPA=4,when i graduate

OpenStudy (anonymous):

GPA =grade point average

OpenStudy (primeralph):

No. To get a 4.0 exactly, you need a 4.0 all through.

OpenStudy (primeralph):

Infinitely, yes due to finite precision. In 8 semesters, no.

OpenStudy (anonymous):

okay thanx thi problem is original ,not from any book i was struggling with it

OpenStudy (primeralph):

You're welcome.

OpenStudy (anonymous):

\[\huge GPA=\frac{\frac{\frac{ x+4}{2}+4}{2}+4}{2}...\] suppose i have many years to study what is the minimum semesters i need to get GPA=4 if my current GPA=3.9

OpenStudy (primeralph):

Depends on precision (number of decimal places). You can never get there if your school uses as many decimal places as possible.

OpenStudy (anonymous):

interesting, thanks @primeralph

OpenStudy (primeralph):

@Jonask If they use 2 decimal places like most schools, then yes you can get there.

OpenStudy (anonymous):

ooh okay but the best way to get it is still 4,my school uses 2 dp

OpenStudy (primeralph):

The best way is 4. If you have a 3.9, you can still get it, but you'll need to work hard.

OpenStudy (whpalmer4):

Your GPA isn't computed by averaging your current GPA with the next grading period's GPA, it is computed by summing the grades over all of the grading periods and then dividing by the number of periods...so the continued fraction you wrote isn't applicable.